OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 4x − 3y = 12 x + 2y = 14 can you show me the steps of doing this

OpenStudy (campbell_st):

label the equations.... (1) and (2) 4x - 3y = 12 (1) x + 2y = 14 (2) The process is to eliminate 1 letter... multiply equation 2 by 4 and then suntract from equation 1 4 x - 3y = 12 - (4x + 8y = 56) - 11y = -44 y = 4 sub into either equation to get x = 6